OpenStudy (anonymous):

Solve the system of equations 5x-y=3 and -10x+2y=-6 Answers: A; 0.-3 B: -3.0 C: 3,12 D: no solution E: infinitely many solutions

OpenStudy (anonymous):

isolate y in the first equation

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

I don't think so... Am I rewriting the first equation?

OpenStudy (anonymous):

ya you want y by itself

OpenStudy (anonymous):

first add y to both sides

OpenStudy (anonymous):

5x=3+y

OpenStudy (anonymous):

then subtract 3 from both sides

OpenStudy (anonymous):

y=5x-3

OpenStudy (anonymous):

now plug in (5x-3) for y in the second equation

OpenStudy (anonymous):

-10x+2(5x-3)=-6

OpenStudy (anonymous):

now try and solve for x

OpenStudy (anonymous):

-10x+4=-6?

OpenStudy (anonymous):

no, foil out 2(5x-3)

OpenStudy (anonymous):

Wouldn't I have to do 2 times 5 then 2 times 3

OpenStudy (anonymous):

2 times 5 plus 2 times -3

OpenStudy (anonymous):

*2 times 5x

OpenStudy (anonymous):

2 times 5x= 10x 2 times =3= -6

OpenStudy (anonymous):

thats right, so we have -10x + 10x - 6 = -6

OpenStudy (anonymous):

simplify that

OpenStudy (anonymous):

Simplify? Like break it down again

OpenStudy (anonymous):

you want to get x by itself to solve it

OpenStudy (anonymous):

Do you minus one of them on both sides

OpenStudy (anonymous):

lets just look at the x's. what is -10x + 10x?

OpenStudy (anonymous):

0..

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

Yayy

OpenStudy (anonymous):

so what should be your answer?

OpenStudy (anonymous):

0,-3?

OpenStudy (anonymous):

no, how did you get -3? anyways, since the x's cancel out there can't be a solution to this system of eqaution

OpenStudy (anonymous):

Ah okay >< sorry, thanks for your help!

OpenStudy (anonymous):

no problem